OpenStudy (anonymous):

I WILL GIVE A MEDAL AND FAN THE PERSON THAT CAN ANSWER THIS. How do you graph f(x)=-2/3x+400? Please explain as well, no need to give me full answer. I want to learn :) PLEASE HELP!!

OpenStudy (anonymous):

its sometimes easier to understand if you switch it to y=mx+b. b=400 and is the y intercept or where the line will sit on the y axis. -2/3 is the slope or m. so on a coordinate plane you would start from the y intercept, or 400, and then plot points going down 2 and right 3. then just conect the point ants and you should be good to go.

OpenStudy (anonymous):

well, i had to do that with y=-2/3x+400. i had to change the equation to function notation and i got f(x)=-2/3x+400. Now its asking me to graph the function, so im guessing i have to do it differently.

OpenStudy (anonymous):

f(x) is the same as y. so they are both graphed the same way. f(x) is just in function notation. both equations should provide you with the same answer.

OpenStudy (anonymous):

I guess that makes sense. Thanks :)

OpenStudy (anonymous):

yup no problem :)

OpenStudy (anonymous):

@TylerElizabeth can you help me with one more?